OpenStudy (anonymous):

Solve the triangle. Round lengths to the nearest tenth and angle measures to the nearest degree. a = 7, b = 7, c = 5

OpenStudy (anonymous):

B. A = 69°, B = 69°, C = 42°??

OpenStudy (anonymous):

sounds good.

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

Try Using cosine law

OpenStudy (anonymous):

\[7^{2} = 7^{2} + 5^{2} - 2(7)(5)cosA\]

OpenStudy (anonymous):

my answers need to be in degrees

OpenStudy (anonymous):

you get \[cosA = \frac{ -25 }{ -70 }\] than do arc cosine. you get angle 69.07

OpenStudy (anonymous):

i need degrees for a b n c ??

OpenStudy (anonymous):

you get angles 41.8 and 69.07. again

OpenStudy (anonymous):

A. A = 70°, B = 70°, C = 40° B. A = 69°, B = 69°, C = 42° C. A = 42°, B = 69°, C = 69° D. A = 69°, B = 42°, C = 69°

OpenStudy (anonymous):

 ?

OpenStudy (anonymous):

Try D.

OpenStudy (anonymous):

have to go.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

deff not 69 deg

OpenStudy (anonymous):

Note that the triangle is isosceles:

OpenStudy (anonymous):

ok ? still confused

OpenStudy (anonymous):

Now use the cosine law:\[\bf a^2=b^2+c^2-2bc \cos(A) \implies 5^2=7^2+7^2-2(7)(7)\cos(A)\]Note that 'A' is the angle opposite side BC. Now re-arrange and solve for cos(A):\[\bf \implies \cos(A)=\frac{-73}{-98}=0.7445 \implies \cos^{-1}(0.7445)=A=41.884 \ degrees\]

OpenStudy (anonymous):

Now because the triangle is isosceles, the remaining two angles B and C are equal so we can find each of them by subtracting A from 180 and dividing the result by 2:\[\bf \angle B=\angle C=\frac{ 180 -41.884}{ 2 }=69.058 degrees\]Rounding off our results to the nearest degree we get:\[\bf \angle A =42 \ degrees\]\[\bf \angle B=\angle C=69 \ degrees\] @britbrat4290

OpenStudy (anonymous):

@britbrat4290 So your answer is correct =]